# Accurate Calibration and Robust LiDAR-Inertial Odometry for Spinning Actuated LiDAR Systems

### v1

@@ Abstract
publicationid: pubid: This article has been accepted for publication in IEEE Robotics and Automation Letters. © 2026 IEEE. Personal use is permitted. All other uses require IEEE permission.
@@ VI Conclusions
Accurate Calibration and Robust LiDAR-Inertial Odometry for Spinning Actuated LiDAR Systems (Supplementary)
@@
S0-A Denavit-Hartenberg Parameters Calibration
@@
The $i$-th plane feature $\pi_{i}$ contains $N_{i}$ laser points. $\mathbf{p}^{L}_{ij}$ represents the $j$-th laser point in the $i$-th plane feature. $\lambda_{\min}(\mathbf{A})$ denotes the minimum eigenvalue of matrix $\mathbf{A}$. The Jacobian matrix $\mathbf{J}$ and the Hessian matrix $\mathbf{H}$ of the cost function (S1) are given by
@@
$\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\mathbf{x}}=\begin{bmatrix}\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{\theta}}&\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{d}}&\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{a}}&\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{\phi}}\end{bmatrix}$. (S3)
@@
$\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{\theta}}=$ $\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\theta_{2}}$ (S4)
@@
$\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{a}}$ $=\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta a_{1}}=\bigg[\mathbf{R}_{z}(\theta_{j})\bigg]_{1},$ (S5)
@@
$\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{a}}$ $=\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta a_{2}}=\bigg[\mathbf{R}_{z}(\theta_{j})\mathbf{R}_{x}(\phi_{1})\mathbf{R}_{z}(\theta_{2})\bigg]_{1},$ (S6)
@@
S0-B Probabilistic Laser Point Representation
@@
$\delta\mathbf{p}^{L}_{j}$ $=\mathbf{A}_{j}\begin{bmatrix}\delta d_{j}\\ \delta\boldsymbol{\omega}_{j}\end{bmatrix}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\mathbf{p}^{L}_{j}}),$ (S7)
@@
$\begin{split}\mathbf{p}_{j}^{M}&=\mathbf{R}_{z}(\theta_{j})(\mathbf{R}_{x}(\phi_{1})\mathbf{R}_{z}(\theta_{2})(\mathbf{R}_{x}(\phi_{2})\mathbf{p}^{L}_{j}+\mathbf{t_{1}})+\mathbf{t}_{2})\\ &=\mathbf{R}_{z}(\theta_{j})(\bar{\mathbf{R}}\mathbf{p}^{L}_{j}+\bar{\mathbf{t}}),\end{split}$ (S8)
@@
$\begin{split}\theta_{j}&=(1-\lambda)\theta_{a}+\lambda\theta_{b}\\ \lambda&=(t_{j}-t_{a})/(t_{b}-t_{a}),t_{a}\leq t_{j}\leq t_{b},\end{split}$ (S9)
@@
$\approx$ $\mathbf{p}_{j}^{M}+\mathbf{J}_{\theta}\delta\theta+\mathbf{J}_{\mathbf{p}^{L}_{j}}\delta\mathbf{p}^{L}_{j},$ (S10)
@@
$\delta\mathbf{p}^{M}_{j}$ $=\begin{bmatrix}\mathbf{J}_{\theta}&\mathbf{J}_{\mathbf{p}^{L}_{j}}\end{bmatrix}\begin{bmatrix}\delta\theta\\ \delta\mathbf{p}^{L}_{j}\end{bmatrix}\sim\mathcal{N}(0,\boldsymbol{\Sigma}_{\mathbf{p^{M}_{j}}}),$ (S11)
@@
$\mathbf{p}^{W}_{j}=\mathbf{R}^{W}_{B}(\mathbf{R}^{B}_{M}\mathbf{p}^{M}_{j}+\mathbf{t}^{B}_{M})+\mathbf{t}^{W}_{B},$ (S12)
@@
$\mathbf{p}^{W,gt}_{j}=$ $\mathbf{R}^{W}_{B}\text{Exp}(\delta\boldsymbol{\theta}^{W}_{B})(\mathbf{R}^{B}_{M}(\mathbf{p}^{M}_{j}+\delta\mathbf{p}^{M}_{j})+\mathbf{t}^{B}_{M})$ (S13)
@@
$\delta\mathbf{p}^{W}_{j}$ $=\mathbf{B}_{j}\begin{bmatrix}\delta\boldsymbol{\theta}^{W}_{B}\\ \delta\mathbf{p}^{M}_{j}\\ \delta\mathbf{t}^{W}_{B}\end{bmatrix}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\mathbf{p}^{W}_{j}}),$ (S14)
@@
S0-C Observability Analysis of LM-Calibr
@@
As illustrated in Fig. S1 and Fig. S2, a smaller minimum eigenvalue indicates reduced observability of LM-Calibr at the associated mounting angle. For spinning omni LiDAR, unobservability occurs only when $\phi_{1}\approx 0$ or $\pi$, with the corresponding mounting configuration shown in Fig. S3(a) and (b). For spinning non-omni LiDAR, LM-Calibr exhibits degeneracy only for the case $\theta_{2}\approx\pm\pi/2$, and the associated mounting angle depicted in Fig. S3(c) and (d). Note that these degenerate mounting configurations are rarely encountered in practice, as they are physically meaningless for increasing FOV. This analysis confirms that LM-Calibr is applicable to various spinning actuated LiDAR mounting configurations.
Figure S1: LM-Calibr observability analysis for spinning omni LiDAR.
Figure S2: LM-Calibr observability analysis for spinning non-omni LiDAR.
Figure S3: The degenerate mounting configurations in LM-Calibr.
S0-D Identifiability Analysis of LM-Calibr
This subsection analyzes the relationship between parameter identifiability in LM-Calibr and geometric conditions through six simulation scenarios. As shown in Fig. S4, the scenarios include isotropic planar distribution (scene_1), x-z planar distribution (scene_2), y-z planar distribution (scene_3), x planar distribution (scene_4), y planar distribution (scene_5), and z planar distribution (scene_6). Each virtual plane measures $10m\times 10m$. The mounting configuration of the spinning Mid360 and spinning Avia are listed in Table S1. The initial values are the ground truth perturbed by $10^{\circ}$ in rotation and $0.1m$ in translation.
Tables S2 and S3 report calibration errors and eigenvalues for each parameter across varying scenarios. Larger eigenvalues imply enhanced parameter identifiability. The results indicate that translational identifiability degrades significantly in single-plane scenarios. This degradation occurs specifically if the plane’s normal vector is parallel to the motor’s spinning axis. Practically, the geometric degeneracy in scene_6 can be circumvented by mounting the sensing system laterally during data acquisition. Therefore, LM-Calibr ensures the identifiability of all calibration parameters via only a single plane, e.g., only ground points.
Figure S4: The six simulation scenarios in the identifiability analysis of LM-Calibr.
TABLE S1: The Mounting Configuration in the Identifiability Analysis of LM-Calibr
@@
TABLE S2: The Calibration Errors ($m$, deg) and Eigenvalues of LM-Calibr in Spinning Mid360 Calibration
@@
TABLE S3: The Calibration Errors ($m$, deg) and Eigenvalues of LM-Calibr in Spinning Avia Calibration
@@
S0-E The Accuracy of EVA-LIO (w/o spin.) in MCD Benchmark
EVA-LIO (w/o spin.) and the baseline methods are evaluated on a representative subset of the MCD Benchmark to demonstrate its generalizability. As shown in Table S4, EVA-LIO (w/o spin.) achieves comparable or superior localization accuracy. This improvement primarily stems from the pre-point uncertainty estimation, as well as from the environmental analysis module, which enables EVA-LIO (w/o spin.) to select appropriate downsampling rates and voxel maps based on the spatial scale.
TABLE S4: Absolute Trajectory Error (m) in MCD Benchmark
@@
S0-F Acceleration Bounds for Stable EVA-LIO Operation
@@
$a||\leq\frac{2\varepsilon_{p}}{T^{2}_{scan}}.$ (S16)
@@
TABLE S5: The System Parameters of Baseline LIOs in CA and CR
@@
TABLE S6: The System Parameters of Baseline LIOs in BG, NTU, KTH, TUHH and Real-world
@@
S0-G System Parameters of Baseline LIOs
To ensure the localization accuracy, specific system parameters of the baseline LIOs are adjusted separately for each sequence. The parameters are listed in Tables S5 and S6.
S0-H Future works

### v2

@@ VI Conclusions
VII Supplementary Material
@@
VII-A Denavit-Hartenberg Parameters Calibration
@@
The $i$-th plane feature $\pi_{i}$ contains $N_{i}$ laser points. $\mathbf{p}^{L}_{ij}$ represents the $j$-th laser point in the $i$-th plane feature. $\lambda_{\min}(\mathbf{A})$ denotes the minimum eigenvalue of matrix $\mathbf{A}$. The Jacobian matrix $\mathbf{J}$ and the Hessian matrix $\mathbf{H}$ of the cost function (6) are given by
@@
$\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\mathbf{x}}=\begin{bmatrix}\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{\theta}}&\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{d}}&\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{a}}&\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{\phi}}\end{bmatrix}$. (8)
@@
$\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{\theta}}=$ $\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\theta_{2}}$ (9)
@@
$\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{a}}$ $=\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta a_{1}}=\bigg[\mathbf{R}_{z}(\theta_{j})\bigg]_{1},$ (10)
@@
$\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta\bar{a}}$ $=\frac{\partial\mathbf{p}^{M}_{ij}(\mathbf{x})}{\partial\delta a_{2}}=\bigg[\mathbf{R}_{z}(\theta_{j})\mathbf{R}_{x}(\phi_{1})\mathbf{R}_{z}(\theta_{2})\bigg]_{1},$ (11)
@@
VII-B Probabilistic Laser Point Representation
@@
$\delta\mathbf{p}^{L}_{j}$ $=\mathbf{A}_{j}\begin{bmatrix}\delta d_{j}\\ \delta\boldsymbol{\omega}_{j}\end{bmatrix}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\mathbf{p}^{L}_{j}}),$ (12)
@@
$\begin{split}\mathbf{p}_{j}^{M}&=\mathbf{R}_{z}(\theta_{j})(\mathbf{R}_{x}(\phi_{1})\mathbf{R}_{z}(\theta_{2})(\mathbf{R}_{x}(\phi_{2})\mathbf{p}^{L}_{j}+\mathbf{t_{1}})+\mathbf{t}_{2})\\ &=\mathbf{R}_{z}(\theta_{j})(\bar{\mathbf{R}}\mathbf{p}^{L}_{j}+\bar{\mathbf{t}}),\end{split}$ (13)
@@
$\begin{split}\theta_{j}&=(1-\lambda)\theta_{a}+\lambda\theta_{b}\\ \lambda&=(t_{j}-t_{a})/(t_{b}-t_{a}),t_{a}\leq t_{j}\leq t_{b},\end{split}$ (14)
@@
$\approx$ $\mathbf{p}_{j}^{M}+\mathbf{J}_{\theta}\delta\theta+\mathbf{J}_{\mathbf{p}^{L}_{j}}\delta\mathbf{p}^{L}_{j},$ (15)
@@
$\delta\mathbf{p}^{M}_{j}$ $=\begin{bmatrix}\mathbf{J}_{\theta}&\mathbf{J}_{\mathbf{p}^{L}_{j}}\end{bmatrix}\begin{bmatrix}\delta\theta\\ \delta\mathbf{p}^{L}_{j}\end{bmatrix}\sim\mathcal{N}(0,\boldsymbol{\Sigma}_{\mathbf{p^{M}_{j}}}),$ (16)
@@
$\mathbf{p}^{W}_{j}=\mathbf{R}^{W}_{B}(\mathbf{R}^{B}_{M}\mathbf{p}^{M}_{j}+\mathbf{t}^{B}_{M})+\mathbf{t}^{W}_{B},$ (17)
@@
$\mathbf{p}^{W,gt}_{j}=$ $\mathbf{R}^{W}_{B}\text{Exp}(\delta\boldsymbol{\theta}^{W}_{B})(\mathbf{R}^{B}_{M}(\mathbf{p}^{M}_{j}+\delta\mathbf{p}^{M}_{j})+\mathbf{t}^{B}_{M})$ (18)
@@
$\delta\mathbf{p}^{W}_{j}$ $=\mathbf{B}_{j}\begin{bmatrix}\delta\boldsymbol{\theta}^{W}_{B}\\ \delta\mathbf{p}^{M}_{j}\\ \delta\mathbf{t}^{W}_{B}\end{bmatrix}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\mathbf{p}^{W}_{j}}),$ (19)
@@
VII-C Observability Analysis of LM-Calibr
@@
As illustrated in Fig. 10 and Fig. 11, a smaller minimum eigenvalue indicates reduced observability of LM-Calibr at the associated mounting angle. For spinning omni LiDAR, unobservability occurs only when $\phi_{1}\approx 0$ or $\pi$, with the corresponding mounting configuration shown in Fig. 12(a) and (b). For spinning non-omni LiDAR, LM-Calibr exhibits degeneracy only for the case $\theta_{2}\approx\pm\pi/2$, and the associated mounting angle depicted in Fig. 12(c) and (d). Note that these degenerate mounting configurations are rarely encountered in practice, as they are physically meaningless for increasing FOV. This analysis confirms that LM-Calibr is applicable to various spinning actuated LiDAR mounting configurations.
Figure 10: LM-Calibr observability analysis for spinning omni LiDAR.
Figure 11: LM-Calibr observability analysis for spinning non-omni LiDAR.
Figure 12: The degenerate mounting configurations in LM-Calibr.
VII-D Identifiability Analysis of LM-Calibr
This subsection analyzes the relationship between parameter identifiability in LM-Calibr and geometric conditions through six simulation scenarios. As shown in Fig. 13, the scenarios include isotropic planar distribution (scene_1), x-z planar distribution (scene_2), y-z planar distribution (scene_3), x planar distribution (scene_4), y planar distribution (scene_5), and z planar distribution (scene_6). Each virtual plane measures $10m\times 10m$. The mounting configuration of the spinning Mid360 and spinning Avia are listed in Table IV. The initial values are the ground truth perturbed by $10^{\circ}$ in rotation and $0.1m$ in translation.
Tables V and VI report calibration errors and eigenvalues for each parameter across varying scenarios. Larger eigenvalues imply enhanced parameter identifiability. The results indicate that translational identifiability degrades significantly in single-plane scenarios. This degradation occurs specifically if the plane’s normal vector is parallel to the motor’s spinning axis. Practically, the geometric degeneracy in scene_6 can be circumvented by mounting the sensing system laterally during data acquisition. Therefore, LM-Calibr ensures the identifiability of all calibration parameters via only a single plane, e.g., only ground points.
Figure 13: The six simulation scenarios in the identifiability analysis of LM-Calibr.
TABLE IV: The Mounting Configuration in the Identifiability Analysis of LM-Calibr
@@ VII-D Identifiability Analysis of LM-Calibr
TABLE V: The Calibration Errors ($m$, deg) and Eigenvalues of LM-Calibr in Spinning Mid360 Calibration
@@
TABLE VI: The Calibration Errors ($m$, deg) and Eigenvalues of LM-Calibr in Spinning Avia Calibration
@@
VII-E The Accuracy of EVA-LIO (w/o spin.) in MCD Benchmark
EVA-LIO (w/o spin.) and the baseline methods are evaluated on a representative subset of the MCD Benchmark to demonstrate its generalizability. As shown in Table VII, EVA-LIO (w/o spin.) achieves comparable or superior localization accuracy. This improvement primarily stems from the pre-point uncertainty estimation, as well as from the environmental analysis module, which enables EVA-LIO (w/o spin.) to select appropriate downsampling rates and voxel maps based on the spatial scale.
TABLE VII: Absolute Trajectory Error (m) in MCD Benchmark
@@
VII-F Acceleration Bounds for Stable EVA-LIO Operation
@@
$a||\leq\frac{2\varepsilon_{p}}{T^{2}_{scan}}.$ (21)
@@
TABLE VIII: The System Parameters of Baseline LIOs in CA and CR
@@
TABLE IX: The System Parameters of Baseline LIOs in BG, NTU, KTH, TUHH and Real-world
@@
VII-G System Parameters of Baseline LIOs
To ensure the localization accuracy, specific system parameters of the baseline LIOs are adjusted separately for each sequence. The parameters are listed in Tables VIII and IX.
VII-H Future works